# Anisotropic Green Coordinates

### v1

@@ 7.1. Cage-based Deformation
We first present the results of anisotropic Green coordinates for cage-based deformation using various symmetric positive definite matrices $\mathbf{A}$. When $\mathbf{A}=\mathbf{I}$, the results are consistent with the isotropic Green coordinates (2008Green). It should be noted that the isotropic method has many desirable properties, such as angle preservation. Thus, one cannot expect that applying an arbitrary matrix $\mathbf{A}$ to induce anisotropic deformation will necessarily yield universally better results than the isotropic formulation. Nevertheless, our method still holds considerable value. This section begins with several representative examples to demonstrate the effects generated by different matrices $\mathbf{A}$, highlighting the diversity of deformation styles and options offered by our method. Additionally, we demonstrate scenarios in which anisotropic deformation potentials yield superior results—whether through enhanced visual quality, tighter adherence to the deformed cage, or reduced isotropic or area distortion. These advantages become more obvious when the input exhibits a high aspect ratio, where appropriate coefficient matrices achieves better outcomes. The formula for computing the distortion can be found in Section C.2.
@@
Figure 4. Cage-based deformation of anisotropic Green coordinate in 2D. Different rotation matrices lead to “stretching” effects along different directions.
@@
We first conduct experiments on 2D images composed of cubic and circular shapes. In Fig. 4, we present the deformation results of a grid-like image under isotropic Green coordinates and anisotropic Green coordinates with eigenvalues of 1.0 and 4.0. The rotation matrices correspond to angles of $\pi/6$ and $\pi/3$; detailed parameter settings can be found in the table of Appendix C.1. We observe that different rotation matrices lead to “stretching” effects along different directions. Fig. 5 illustrates another example, in which the eigenvalues of the anisotropic coefficient matrix in the non-diagonal matrix are typically 1.0 and 2.0. The results demonstrate that isotropic Green coordinates are angle preserving and map infinitesimal circles to infinitesimal circles. However, the resulting effect are fixed, which also limits the flexibility and diversity of deformation. Our method enables a wide variety of deformation styles, and appropriate parameter settings can achieve desired effects. For instance, the leftmost image in the middle row of Fig. 5 minimizes area distortion.
@@
Figure 5. Cage-based deformation results of an image containing circles and cubes.
Fig. 6, Fig. 7, and Fig. 8 show the cage-based deformation results of a fish, an elephant, and a giraffe model, respectively. It can be observed that different matrices control the “fatness” or “thinness” of the deformed objects and produce diverse deformation effects. The crab example in Fig. 9 further demonstrates that appropriate anisotropic settings can improve the deformation effect, exhibiting better consistency with the cage and resulting in more natural deformations compared to the input. Moreover, the anisotropic results exhibit lower isometric distortion (0.0245 vs. 0.0328) and area distortion (0.1340 vs. 0.1829).
@@
Figure 6. Cage-based deformation results of a fish model.
Figure 7. Cage-based deformation results of an elephant model.
Figure 8. Cage-based deformation results of a giraffe model.
@@
Figure 9. Cage-based deformation results of a crab model.
@@
Figure 10. Cage-based deformation results of a bar model in 3D.
@@
Figure 11. Cage-based deformation results of a bench model in 3D.
Figure 12. Cage-based deformation results of a skybox model in 3D.
@@
Next, we move on to the results of 3D deformation. We first present the deformation of a bar model in Fig. 10, where the coordinate axes are indicated in the lower-left corner of the image. The deformation is mainly performed along the y-axis, while the long side is oriented along the z-axis. We show the anisotropic results aligned with both y-axis and z-axis. It can be observed that different matrices produce results with varying degrees of sharpness. Fig. 11 shows the results of a bench model, where different anisotropic matrices control the degree of “upward bending” in the chair’s crossbeam. In the skybox model of Fig. 12, the first row shows results for diagonal matrices, while the second row presents outcomes with non-diagonal matrices, specifically $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,0)$, $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,\pi/4)$ and $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\frac{\pi}{3},\frac{\pi}{4},\frac{\pi}{6})$ from left to right. The experimental results further illustrate the wide range of deformation effects achieved by our approach.
@@ 7.2. Variational Shape Deformation
Here, we conduct experiments on variational shape deformation using different coefficient matrices. The method is reimplemented based on the technique proposed by (2009VariationalGC) using Eq. (54). For the same input, all parameters are held constant except for the anisotropic matrix. Fig. 13 presents the results of variational shape deformation in 2D. The user-specified pre-deformation positions are indicated by red dots, while the target positions are marked with blue dots. The input cage is illustrated using orange-yellow lines. The parameters for the 2D deformation are configured as $\lambda_{1}=100,\lambda_{2}=10$ and $\lambda_{3}=0.1$. Additionally, 10 boundary points are sampled to compute the Hessian near each face for the first two inputs, while 2 are sampled near each face for the last input. The results show that varying the anisotropic matrices leads to different deformation outcomes, thereby broadening the range of deformation effects. Fig. 14 further shows the results of 3D variational deformation, where red and blue points are used to mark constraint points. We also manually annotate the constraint points that are occluded during rendering. For the Bar model, the parameters are set to $\lambda_{1}=1000,\lambda_{2}=0.5$ and $\lambda_{3}=0.001$, while for the Botijo model, the parameters are $\lambda_{1}=100,\lambda_{2}=0.05$ and $\lambda_{3}=1.0$. The results further illustrate the different effects produced by anisotropic deformations.
Figure 14. Results of variational shape deformation in 3D.
@@ 8. Conclusion
In this work, we introduce anisotropic Green coordinates, which incorporate a symmetric positive definite coefficient matrix within the divergence operator to enable diverse deformation styles in both cage-based deformation and variational shape deformation. We derive an integral identity formulation for the anisotropic Laplacian equation, ensuring that our method satisfies both linear reproduction and translation invariance. Moreover, anisotropic Green coordinates, along with their gradients and Hessians, admit closed-form expressions, enabling efficient deformation computation and supporting variational shape deformation with an as-rigid-as-possible energy term while accommodating user-specified positional constraints.
Our method still has certain limitations. The primary one is that $\mathbf{A}$ remains globally invariant and has not been optimized for deformation constraints. Additionally, Green coordinates cannot simultaneously satisfy both Dirichlet and Neumann boundary conditions. We adopt Green coordinates as the foundation of our approach because of their theoretical elegance and strong compatibility with our concept of anisotropic deformation. It is our hope that this idea offers novel insights for anisotropic cage-based and variational deformation, and that this work inspires further research in the field.
See pages - of img/Supplementary.pdf

### v2

@@ 7.1. Cage-based Deformation
First, we present the results of cage-based deformation using anisotropic Green coordinates with various control matrices. When $\mathbf{A}=\mathbf{I}$, our results align with those of the traditional isotropic technique (Lipman et al., 2008). The isotropic method ensures strict angle preservation in 2D, whereas anisotropy introduces some degree of shearing. However, we will demonstrate that this capability holds considerable practical value. We start with several representative examples to demonstrate the effects generated by different matrices $\mathbf{A}$, highlighting the diversity of deformation styles and options offered by our method. Additionally, we demonstrate scenarios in which anisotropic deformation potentials yield superior results,characterized by enhanced visual quality, tighter adherence to the deformed cage, or reduced isotropic or area distortion. These improvements are particularly evident in inputs with a high aspect ratio. Please refer to Appendix C.2 for the formula used to compute distortion. Subsequently, we will compare our method with mean value coordinates (Floater, 2003) and harmonic coordinates (Joshi et al., 2007).
@@
Figure 4. Cage-based deformation of anisotropic Green coordinate in 2D. Different rotation matrices lead to varying “stretching” effects along different directions.
@@
We first conduct experiments on 2D images composed of cubic and circular shapes. In Fig. 4, we present the deformation results of a grid-like image. The eigenvalues of the matrix are $1.0$ and $4.0$, while the rotation angles are $\pi/6$ and $\pi/3$. Detailed parameter settings can be found in the table of Appendix C.1. The schematic coordinate frames associated with $\lambda_{1}$, $\lambda_{2}$ and $\theta$ are also indicated in the figure. We observe that different rotation matrices lead to varying degrees of “stretching” along different directions, which intuitively demonstrates the effects of anisotropy. Fig. 5 illustrates another example involving basic shapes such as cubes and circles. The eigenvalues of the non-diagonal matrix are typically $1.0$ and $2.0$. While isotropic Green coordinates are angle-preserving and map infinitesimal circles to infinitesimal circles, their deformation effect is fixed, which limits its flexibility and diversity. In contrast, our method enables a wide variety of deformation styles. Furthermore, regarding area distortion, the leftmost image in the middle row yields the lowest value (0.420) compared to the isotropic result (0.595).
@@
Figure 5. Cage-based deformation results of an image containing circles and cubes. Anisotropic Green coordinates brings diverse deformation effects.
@@
Figure 6. Cage-based deformation results of a fish, an elephant, and a giraffe model, respectively.
@@
Figure 7. Cage-based deformation results of a crab model. The tuple $(a,b)$ represents the isometric distortion and area distortion, respectively.
@@
Figure 8. Comparison of cage-based deformation results using our method, MVC (Floater, 2003), and HC (Joshi et al., 2007).
@@
Figure 9. Cage-based deformation results of a bar, a bench and a skybox model in 3D, respectively.
Fig. 6 shows the cage-based deformation results of a fish, an elephant, and a giraffe model, respectively. It can be observed that different matrices control the “fatness” or “thinness” of the deformed objects, thereby affording artists and animators a more versatile range of deformation choices. The crab example in Fig. 7 further demonstrates that appropriate anisotropic settings can improve deformation effects. In the figure, the tuple $(a,b)$ represents the isometric distortion and area distortion, respectively. We can observe that the anisotropic results yield more natural deformations and exhibit lower isometric and area distortion.
To further demonstrate the advantages of our method, we compare our method with mean value coordinates (MVC) (Floater, 2003) and harmonic coordinates (HC) (Joshi et al., 2007) in Fig. 8. It is evident that MVC and HC lead to noticeable shearing and unnatural deformations. Our method yields different deformation results compared to the isometric Green coordinates. Meanwhile, the quasi-conformal property of our method also guarantees a certain degree of shape preservation.
Figure 11. Results of variational shape deformation in 3D. Red and blue dots mark the spatial constraints.
@@
We then present the results of 3D deformation. In Fig. 9, we show the results of a bar, a bench, and a skybox model, respectively. We also mark the spatial coordinate axes in the figure. In the bar model, the deformation is mainly observed along the y-axis, while the long side is oriented along the z-axis. We show the anisotropic deformation results along the y-axis or z-axis in the coefficient matrix. It can be observed that different matrices produce deformation effects with varying degrees of sharpness. In the bench model, we observe that different anisotropic matrices control the degree of upward bending in the chair’s crossbeam. In the skybox model, the first row displays results obtained using diagonal matrices, while the second row presents outcomes for non-diagonal matrices. Specifically, from left to right, the matrices are $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,0)$, $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,\pi/4)$ and $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\frac{\pi}{3},\frac{\pi}{4},\frac{\pi}{6})$. These experimental results further demonstrate the wide range of deformation effects achieved by our approach.
@@ 7.2. Variational Shape Deformation
In this subsection, we conduct experiments on variational shape deformation proposed in Section 6. Specifically, we establish the deformation mapping based on anisotropic Green coordinates as Eq. (47), and optimize $\mathbf{a}$ and $\mathbf{b}$ by enforcing spatial constraints and as-rigid-as-possible constraints, as formulated in Eq. (55). The spatial constraints dictate the deformation of $\mathbf{q}_{i}$ to $\mathbf{f}_{i}$ for i=1,2,,ri=1,2,...,r. The as-rigid-as-possible constraints aim to preserve rigidity during deformation as much as possible for some selected positions. Similar to previous methods, we sample a set of points $\mathbf{m}_{i}$ for i=1,2,,di=1,2,...,d along the medial axis. The Hessian energy term and the regularization term ensure the smoothness of the deformation and the stability of optimization, respectively. Since harmonic functions satisfy the extremum principle at the boundary, this property is inherited by our approach, in the sense that boundary points map to boundary points under linear transformation. Consequently, we sample the boundary points $\mathbf{w}_{i}(i=1,2,...,k)$ near the cage faces.
Figures 10 and 11 illustrate the 2D and 3D results, respectively, which are obtained using the same input and keeping all parameters constant except for the anisotropic matrix. Red dots denote the user-specified pre-deformation positions, while blue dots mark the target positions. The 2D input cages are illustrated in orange lines. The parameters are configured as $\lambda_{1}=100,\lambda_{2}=10$ and $\lambda_{3}=0.1$. In the first two examples, we sample ten boundary points near each cage face to compute the Hessians, whereas for the final example, we sample two points near each face. The results demonstrate that varying the anisotropic matrices yields different deformation outcomes, thereby expanding the range of achievable deformation effects. In the 3D scenario, we also use red and blue points to mark the points of spatial constraints. Furthermore, we manually annotate the constraints that are occluded during rendering. The initial cages are dense, subdivided rectangular triangle meshes enclosing the entire input. We sample two points near each face to compute the Hessians. The parameters are set to $\lambda_{1}=1000,\lambda_{2}=0.5$ and $\lambda_{3}=0.001$ for the bar model (left), and to $\lambda_{1}=100,\lambda_{2}=0.05$ and $\lambda_{3}=0.001$ for the Botijo model (right). These results further demonstrate the variety of deformation effects achieved by our method.
@@ 8. Conclusion
In this work, we introduce anisotropic Green coordinates, which incorporate a symmetric positive definite coefficient matrix into the divergence operator of the Laplacian equation. This approach enables a variety of deformation styles for both cage-based deformation and variational shape deformation. We derive the Green’s identity based on the anisotropic Laplacian equation and demonstrate that our method satisfies both linear precision and translation invariance. Furthermore, anisotropic Green coordinates, along with their gradients and Hessians, admit closed-form expressions. This facilitates the efficient computation of deformation mappings and supports variational shape deformation using as-rigid-as-possible energy, while accommodating user-specified positional constraints.
The primary limitation of our method is that the matrix $\mathbf{A}$ remains globally invariant and is not optimized based on deformation constraints, which we leave to future work. Additionally, Green coordinates cannot satisfy both Dirichlet and Neumann boundary conditions simultaneously. We adopt Green coordinates as the foundation of our approach because of their theoretical elegance and strong compatibility with the concept of anisotropy. We anticipate that this framework will offer novel insights for anisotropic space deformation and inspire further research in this field.